If a 1000g sample of Au-198 is held for a little over 5.4 days, how much of the sample will remain?


1
Expert's answer
2021-02-01T03:47:14-0500

"\dfrac{5.4days}{2.7days} =" 2 half-lifes


"\implies" "(\dfrac{1}{2})^2 = \dfrac{1}{4}"


"\dfrac{1}{4}x1000g = 250g" will remain